In series circuit, the voltmeter reads 120V. Find what is (a.) the combined resistance (b) the potential differences across the combination (c) the voltage across the 30ꭥ resistor and (d) the current flows through the 10ꭥ resistor?


1
Expert's answer
2022-03-17T09:36:54-0400

Explanations & Calculations


  • If the resistors here in series are those 30 ohms and 10 ohms.
  • Then the combined resistance is 30+10 = 40 Ohms.
  • The potential difference is what the voltmeter measures and it could be 120V according to the diagram you have.
  • If the voltage 120V is across the two resistors, then the potential difference across 30 Ohm is

"\qquad\qquad
\begin{aligned}
\small V_{30}&=\small 120\times\frac{30}{30+10}=90\,V
\end{aligned}"

  • Since the current through each resistor is the equivalent current that flows in the circuit,

"\qquad\qquad
\begin{aligned}
\small I_{eq}&=\small \frac{120\,V}{40\,\Omega}\\
&=\small 3\,A 
\end{aligned}"